1. How much work is required to move a charge of 4nC from a point 2m away to a


point 0.5 m away from a point charge of 60nC? What is the potential difference


between these points?



2. An electron is accelerated from rest through a potential difference of 15V. What is the change in electric potential energy of the electron?



3. Two-point charges Q1 = 4.5 μC and Q2 = 8.5 μC are initially very far apart. They are then brought together, with a final separation of 3.5 mm. How much work does it take to bring them together?



4. A 0.500 cm diameter plastic sphere, used in a static electricity demonstration, has a uniformly distributed 40.0 pC charge on its surface. What is the potential near its surface?

1
Expert's answer
2022-02-20T08:20:25-0500

1.

"W=kq_1q_2(1/r_2-1/r_1)\\
=9*10^9*4*10^{-9}*60*10^{-9}(1/0.5-1/2)\\
=3.24*10^{-6}\:\rm J"

2.

"\Delta E_p=qV=1.6*10^{-19}\: \rm C*15\: \rm V=2.4*10^{-18}\: J"

3.

"W=kq_1q_2/r\\
=9*10^9*4.5*10^{-6}*8.5*10^{-6}/0.0035\\
=98\:\rm J"

4.

"V=kq/r\\
=9*10^9*40.0*10^{-12}/0.05=7.2\:\rm V"